OpenStudy (anonymous):

whats is g(4)if g(x)=x^2-x

OpenStudy (striker1298):

hey jose i can help

OpenStudy (anonymous):

help please

OpenStudy (striker1298):

plug in 4 for every x: \[g(4)=(4)^2-4\]

OpenStudy (striker1298):

You then get: \[g(4)=16-4\] Which then equals: \[g(4)=12\]